# PAC in DESI. I. Galaxy Stellar Mass Function into the $10^{6}{\rm M}_{\odot}$ Frontier

### v2

@@ 1 Introduction
Over the past two decades, both cosmological spectroscopic surveys (York et al., 2000; Colless et al., 2001; Jones et al., 2004; Le Fèvre et al., 2005; Driver et al., 2011; Newman et al., 2013; Garilli et al., 2014) and photometric surveys (York et al., 2000; The Dark Energy Survey Collaboration, 2005; Heymans et al., 2012; de Jong et al., 2015; Aihara et al., 2018a; Dey et al., 2019) have significantly contributed to shaping our understanding of galaxy formation and cosmology. These two types of surveys possess unique advantages and disadvantages. Spectroscopic surveys can offer precise redshift measurements for studying 3D galaxy clustering and assessing galaxy properties. However, they still focus on bright and selected tracers, leading to complex selection effects. Photometric surveys are significantly deeper and more complete than spectroscopic surveys, and the gap is expected to widen further in the future. Photometric surveys can also provide spatially resolved galaxy properties such as morphology and shape. Nonetheless, obtaining precise redshift measurements from photometric surveys remains a challenging task (Newman & Gruen, 2022). Moreover, due to their distinct characteristics, spectroscopic and photometric surveys are often aligned with different scientific objectives. For instance, photometric surveys are typically driven by weak lensing measurements (Kaiser & Squires, 1993; Bartelmann & Schneider, 2001), whereas spectroscopic surveys are primarily pursued to study baryon acoustic oscillations (BAO; Seo & Eisenstein, 2003) and redshift space distortions (RSD; Kaiser, 1987).
@@
If we can integrate these two types of surveys and make full use of their respective strengths, we can generate a wealth of information to advance our understanding of galaxy formation and cosmology. To fully leverage the deep photometric surveys, as both photometric and spectroscopic objects trace the comic web, it is theoretically possible to statistically assign redshifts by cross-correlating them with spectroscopic tracers, eliminating the use of photometric redshift (photoz). However, in practice, a well-thought-out design is essential to achieve this goal. Various existing approaches, such as the clustering-z method in weak lensing (Newman, 2008; Ménard et al., 2013), have made attempts in this direction, but they still rely on certain assumptions, such as those related to galaxy bias (Schmidt et al., 2013; Rahman et al., 2015).
@@
To achieve this, building upon the work of Wang et al. (2011), we introduced the Photometric objects Around Cosmic webs (PAC) method in Xu et al. (2022a, hereafter Paper I). The PAC method estimates the excess surface density, $\bar{n}_{2}w_{{\rm{p}}}$, of photometric objects with specific physical properties, such as stellar mass, luminosity, rest-frame colour, and star formation rate, around spectroscopic tracers. Utilizing Stage-III cosmological spectroscopic and photometric surveys such as the Sloan Digital Sky Survey (SDSS) (York et al., 2000; Dawson et al., 2013), the Hyper Suprime-Cam (HSC) SSP Survey (Aihara et al., 2018a), and the DECam Legacy Survey (DECaLS, Dey et al., 2019), we have conducted a series of investigations using the PAC method. These studies have yielded innovative measurements and results in various aspects of galaxy formation, including precise measurements of the galaxy stellar mass function (GSMF; Xu et al., 2022b, hereafter Paper III) and the stellar mass-halo mass relation (SHMR; Xu et al., 2023a, hereafter Paper IV) down to stellar mass $10^{8.0}{\rm M}_{\sun}$, as well as their evolution up to redshift 0.7. Additionally, we have investigated galaxy assembly bias (Xu & Jing, 2022), environmental quenching (Zheng et al., 2024; Zheng et al., 2025), and the quasar-halo connection (Gui et al., 2024). In Paper IV, we also developed accurate galaxy-halo connections for the SDSS spectroscopic CMASS and LOWZ samples, which have been applied to study the intrinsic alignment of massive galaxies (Xu et al., 2023b) and gravitational lensing magnification around them (Xu et al., 2024). Furthermore, many other studies utilized a concept similar to the PAC method, with a focus on studying the distribution of satellite galaxies (Hansen et al., 2009; Guo et al., 2011; Jiang et al., 2012; Wang & White, 2012; Guo et al., 2012; Wang et al., 2014a; Lan et al., 2016; Wang et al., 2021; Tinker et al., 2021; Wang et al., 2025). These studies underscore the well-established status of the PAC method and have already contributed to enhancing our understanding of galaxy formation.
@@
We are now entering the Stage-IV era, with various next-generation cosmological spectroscopic and photometric surveys either commencing or planning to release their data in the next few years, including the Dark Energy Spectroscopic Instrument (DESI; DESI Collaboration et al., 2016a), the Subaru Prime Focus Spectrograph (PFS; Takada et al., 2014), the Legacy Survey of Space and Time (LSST; Ivezić et al., 2019), Euclid (Laureijs et al., 2011), the Chinese Space Station Optical Survey (CSS-OS; Gong et al., 2019), and surveys conducted with the Nancy Grace Roman Space Telescope (Roman; Spergel et al., 2015). DESI aims to provide more than 40 million spectroscopic tracers up to $z\sim 3.5$, which is approximately 40 times greater than the number available in SDSS. LSST, Euclid, CSS-OS and Roman will provide large area multi-band galaxy images ranging from near-ultraviolet to near-infrared to unparalleled depth, reaching approximately $r\sim 27$ or even deeper, which is $2-3$ magnitudes deeper than DECaLS we utilized in previous studies. The Stage-IV surveys present promising opportunities for us to employ the PAC method in exploring fainter sources and earlier cosmic epochs and obtaining more precise measurements. These innovative measurements can contribute to addressing fundamental questions regarding the nature of dark matter and dark energy, the epoch of reionization, structure formation in the universe, and the formation of galaxies.
@@
In this paper, as the first application of the PAC method in the Stage-IV era, we focus on measuring the local GSMF down to much lower mass limits than previously achieved. The GSMF is a crucial measurement for understanding galaxy evolution. The amplitude and shape of the GSMF encode vital information about galaxy star formation and quenching mechanisms. Consequently, it has been utilized to infer and constrain galaxy formation models in numerous studies ranging from simple empirical models to numerical simulations (Peng et al., 2010; Behroozi et al., 2013; Schaye et al., 2015; Henriques et al., 2015; Pillepich et al., 2018; Davé et al., 2019; Schaye et al., 2023). In previous works, large flux-limited spectroscopic surveys, particularly the Two Degree Field Galaxy Spectroscopic Survey (2dF-GRS; Colless et al., 2001), SDSS, and Galaxy and Mass Assembly (GAMA; Driver et al., 2011), are commonly employed to investigate the local GSMF (Cole et al., 2001; Baldry et al., 2008; Li & White, 2009; Driver et al., 2022). However, due to depth limitations, accurate and reliable measurements of the local GSMF from spectroscopic surveys remain constrained to relatively high masses. At lower masses, measurements are typically refined to the Local Volume, which lies within or around the Local Void (Tully et al., 2008; Peebles & Nusser, 2010; McConnachie, 2012; Karachentsev & Makarova, 2019; Böhringer et al., 2020), often leading to underestimation of the GSMF. Using the ELUCID simulation (Wang et al., 2014b, 2016b), which was run based on the reconstructed initial density field of the nearby universe, Chen et al. (2019) quantified the underestimation of the GSMF by galaxy surveys with various magnitude limits. They identified significant underestimation at the low-mass end in current surveys, starting at stellar mass $10^{8.0}{\rm M}_{\sun}-10^{9.0}{\rm M}_{\sun}$. Therefore, measuring the low-mass end of the GSMF using spectroscopic surveys alone remains highly challenging.
@@
To reach lower stellar masses, we employ the PAC method in combination with deeper photometric surveys, rather than relying solely on spectroscopic data, as have been achieved in Paper III. Notably, comparing to the stellar mass limits of $10^{8.2}{\rm M}_{\sun}$ reached in Paper III, we enhance various aspects of the previous PAC method and take a significant step into the $10^{6.0}{\rm M}_{\sun}$ frontier with the DESI Y1 Bright Galaxy Survey (BGS) Bright sample and DECaLS photometric sample. This advance is facilitated by leveraging numerous low-redshift tracers available in BGS, enabling accurate $\bar{n}_{2}w_{{\rm{p}}}$ measurements at lower redshifts. Furthermore, we derive the GSMF for blue and red galaxy samples separately, offering deeper insights into the galaxy population. Our results mark significant progress in the measurement of the local GSMF and have the potential to impact our understanding of dark matter and reionization. Specifically, the number density of galaxies with stellar masses around $10^{6.0}{\rm M}_{\sun}$ becomes sensitive to dark matter particle mass (Colín et al., 2000; Bode et al., 2001; Viel et al., 2005; Lovell et al., 2012; Macciò et al., 2013; Horiuchi et al., 2016; Bose et al., 2017) and to the epoch and physics of reionization (Efstathiou, 1992; Loeb & Barkana, 2001; Okamoto et al., 2008; Benítez-Llambay et al., 2017; Bose et al., 2018). Additionally, our findings can also contribute to advancing the understanding of dwarf galaxy formation.
@@
We introduce the DESI and DECaLS data in Section 2. Section 3 provides a detailed description of the PAC method, including various improvements made to enhance it. The measurements and the resulting GSMF are presented in Section 4. In Section 5, we provide a concise summary. We adopt the cosmology with $\Omega_{\rm m}=0.268$, $\Omega_{\Lambda}=0.732$ and $H_{0}=71{\rm\ km/s/Mpc}$ throughout the paper (Hinshaw et al., 2013).
@@ 2.1 DESI Y1 BGS
DESI is a prominent Stage-IV dark energy survey with the goal of obtaining spectra for approximately 40 million extragalactic objects over a five-year period (Levi et al., 2013; DESI Collaboration et al., 2016a, b, 2022, 2024a, 2024b). Covering more than 14,000 $\mathrm{deg^{2}}$, the survey is conducted using a multi-object fibre-fed spectrograph mounted on the prime focus panel of the 4-meter Mayall Telescope at Kitt Peak National Observatory (DESI Collaboration et al., 2022). The spectrometer operates in the wavelength range of $3600-9800\,$Å and can allocate fibres to 5,000 objects during a visit (DESI Collaboration et al., 2016b; Silber et al., 2023; Miller et al., 2024; Poppett et al., 2024). Additional pipelines supporting the DESI experiment are detailed in Guy et al. (2023); Bailey et al. (2025); Raichoor et al. (2025); Schlafly et al. (2023); Myers et al. (2023). The parent catalogue used for DESI target selections is constructed from Data Release 9 of the DESI Legacy Imaging Surveys (Zou et al., 2017; Dey et al., 2019). The photometric data contains three optical bands $grz$ from the DECam Legacy Survey (DECaLS; Dey et al., 2019), the Dark Energy Survey (DES; The Dark Energy Survey Collaboration, 2005), the Beijing-Arizona Sky Survey (BASS; Zou et al., 2017) and the Mayall $z$-band Legacy Survey (MzLS).
As part of its core observations, DESI is conducting the Bright Galaxy Survey (BGS; Hahn et al., 2023). BGS spans 14,000 $\mathrm{deg^{2}}$ footprint and includes low-redshift galaxies with $z<0.6$ that can be observed during bright time, when the night sky is approximately 2.5×2.5$\times$2.5 × brighter than nominal dark conditions. BGS provides two galaxy samples: the BGS Bright sample, a $r<19.5$ magnitude-limited sample of approximately 10 million galaxies, and the BGS Faint sample, a fainter $20.17519.5<r<$ 20.175 sample of approximately 5 million galaxies selected based on surface brightness and colour. The selection and completeness of the BGS samples are detailed in Hahn et al. (2023).
@@
where $w_{\rm{comp}}$ corrects for fibre-assignment incompleteness, and $w_{\rm zfail}$ adjusts for changes in the relative redshift success rates (Ross et al., 2025).
In Figure 1, we present the redshift and stellar mass distributions of the BGS Bright sample within the DECaLS region. We compute their physical properties by applying the SED fitting code CIGALE (Boquien et al., 2019), utilizing the $grz$ band fluxes from DECaLS and redshifts from DESI spectra. The calculations are based on the stellar population synthesis models by Bruzual & Charlot (2003), which employ a Chabrier (2003) initial mass function and a star formation history characterized by a delayed exponential function $\phi(t)\approx t\exp{(-t/\tau)}$. We consider three metallicities, specifically $Z/Z_{\odot}=0.4,\ 1,\ \text{and}\ 2.5$, where $Z_{\odot}$ represents the metallicity of the Sun. Additionally, we incorporate the Calzetti et al. (2000) extinction law to account for dust reddening, within the range of $0.50<E(B-V)<$ 0.5.
Since we aim to study the blue and red populations separately, Figure 2 presents the rest-frame $g-r$ colour vs. stellar mass distributions, weighted by 1/$V_{\rm{max}}$, where $V_{\rm{max}}$ corresponds to $z_{\rm{max}}$, the maximum redshift at which a galaxy can still meet our sample selection criteria. The $V_{\rm{max}}$ values are taken from Wang et al. (2024). We apply the colour cut $g-r=0.04\log_{10}(M_{*}/{\rm M}_{\sun})+0.2$ to separate the blue and red populations. This cut is chosen to closely match the one used in Zheng et al. (2024), where they studied environmental quenching using PAC and SDSS data. Their colour cut, based on $u-r$, could not be directly used because DECaLS does not include $u$-band measurements. Instead, we used $g-r$ and determined the cut based on the relationship between the $u$ and $g$ bands from SDSS data.
To calculate the correlation functions, we use a random catalogue provided by DESI (Ross et al., 2025), which matches the footprint and redshift distribution of the BGS sample used in our analysis.

### v3

@@ 1 Introduction
Over the past two decades, both cosmological spectroscopic surveys (2000AJ....120.1579Y; 2001MNRAS.328.1039C; 2004MNRAS.355..747J; 2005A&A...439..845L; 2011MNRAS.413..971D; 2013ApJS..208....5N; 2014A&A...562A..23G) and photometric surveys (2000AJ....120.1579Y; 2005astro.ph.10346T; 2012MNRAS.427..146H; 2015A&A...582A..62D; 2018PASJ...70S...4A; 2019AJ....157..168D) have significantly contributed to shaping our understanding of galaxy formation and cosmology. These two types of surveys possess unique advantages and disadvantages. Spectroscopic surveys can offer precise redshift measurements for studying 3D galaxy clustering and assessing galaxy properties. However, they still focus on bright and selected tracers, leading to complex selection effects. Photometric surveys are significantly deeper and more complete than spectroscopic surveys, and the gap is expected to widen further in the future. Photometric surveys can also provide spatially resolved galaxy properties such as morphology and shape. Nonetheless, obtaining precise redshift measurements from photometric surveys remains a challenging task (2022ARA&A..60..363N). Moreover, due to their distinct characteristics, spectroscopic and photometric surveys are often aligned with different scientific objectives. For instance, photometric surveys are typically driven by weak lensing measurements (1993ApJ...404..441K; 2001PhR...340..291B), whereas spectroscopic surveys are primarily pursued to study baryon acoustic oscillations (BAO; 2003ApJ...598..720S) and redshift space distortions (RSD; 1987MNRAS.227....1K).
@@
If we can integrate these two types of surveys and make full use of their respective strengths, we can generate a wealth of information to advance our understanding of galaxy formation and cosmology. To fully leverage the deep photometric surveys, as both photometric and spectroscopic objects trace the comic web, it is theoretically possible to statistically assign redshifts by cross-correlating them with spectroscopic tracers, eliminating the use of photometric redshift (photoz). However, in practice, a well-thought-out design is essential to achieve this goal. Various existing approaches, such as the clustering-z method in weak lensing (2008ApJ...684...88N; 2013arXiv1303.4722M), have made attempts in this direction, but they still rely on certain assumptions, such as those related to galaxy bias (2013MNRAS.431.3307S; 2015MNRAS.447.3500R).
@@
To achieve this, building upon the work of 2011ApJ...734...88W, we introduced the Photometric objects Around Cosmic webs (PAC) method in 2022ApJ...925...31X. The PAC method estimates the excess surface density, $\bar{n}_{2}w_{{\rm{p}}}$, of photometric objects with specific physical properties, such as stellar mass, luminosity, rest-frame colour, and star formation rate, around spectroscopic tracers. Utilizing Stage-III cosmological spectroscopic and photometric surveys such as the Sloan Digital Sky Survey (SDSS) (2000AJ....120.1579Y; 2013AJ....145...10D), the Hyper Suprime-Cam (HSC) SSP Survey (2018PASJ...70S...4A), and the DECam Legacy Survey (DECaLS, 2019AJ....157..168D), we have conducted a series of investigations using the PAC method. These studies have yielded innovative measurements and results in various aspects of galaxy formation, including precise measurements of the galaxy stellar mass function (GSMF; 2022ApJ...939..104X, hereafter 2022ApJ...939..104X) and the stellar mass-halo mass relation (SHMR; 2023ApJ...944..200X, hereafter 2023ApJ...944..200X) down to stellar mass $10^{8.0}{\rm M}_{\sun}$, as well as their evolution up to redshift 0.7. Additionally, we have investigated galaxy assembly bias (2022ApJ...926..130X), environmental quenching (2024ApJ...969..129Z; 2025arXiv250100986Z), and the quasar-halo connection (2024ApJ...967...17G). In 2023ApJ...944..200X, we also developed accurate galaxy-halo connections for the SDSS spectroscopic CMASS and LOWZ samples, which have been applied to study the intrinsic alignment of massive galaxies (2023ApJ...954....2X) and gravitational lensing magnification around them (2024ApJ...973..102X). Furthermore, many other studies utilized a concept similar to the PAC method, with a focus on studying the distribution of satellite galaxies (2009ApJ...699.1333H; 2011MNRAS.417..370G; 2012ApJ...760...16J; 2012MNRAS.424.2574W; 2012MNRAS.427..428G; 2014MNRAS.442.1363W; 2016MNRAS.459.3998L; 2021ApJ...919...25W; 2021MNRAS.505.5370T; 2025arXiv250303317W). These studies underscore the well-established status of the PAC method and have already contributed to enhancing our understanding of galaxy formation.
@@
We are now entering the Stage-IV era, with various next-generation cosmological spectroscopic and photometric surveys either commencing or planning to release their data in the next few years, including the Dark Energy Spectroscopic Instrument (DESI; 2016arXiv161100036D), the Subaru Prime Focus Spectrograph (PFS; 2014PASJ...66R...1T), the Legacy Survey of Space and Time (LSST; 2019ApJ...873.. [^1] I), Euclid (2011arXiv1110.3193L), the Chinese Space Station Optical Survey (CSS-OS; 2019ApJ...883..203G), and surveys conducted with the Nancy Grace Roman Space Telescope (Roman; 2015arXiv150303757S). DESI aims to provide more than 40 million spectroscopic tracers up to $z\sim 3.5$, which is approximately 40 times greater than the number available in SDSS. LSST, Euclid, CSS-OS and Roman will provide large area multi-band galaxy images ranging from near-ultraviolet to near-infrared to unparalleled depth, reaching approximately $r\sim 27$ or even deeper, which is 232-3 magnitudes deeper than DECaLS we utilized in previous studies. The Stage-IV surveys present promising opportunities for us to employ the PAC method in exploring fainter sources and earlier cosmic epochs and obtaining more precise measurements. These innovative measurements can contribute to addressing fundamental questions regarding the nature of dark matter and dark energy, the epoch of reionization, structure formation in the universe, and the formation of galaxies.
@@
In this paper, as the first application of the PAC method in the Stage-IV era, we focus on measuring the local GSMF down to much lower mass limits than previously achieved. The GSMF is a crucial measurement for understanding galaxy evolution. The amplitude and shape of the GSMF encode vital information about galaxy star formation and quenching mechanisms. Consequently, it has been utilized to infer and constrain galaxy formation models in numerous studies ranging from simple empirical models to numerical simulations (2010ApJ...721..193P; 2013ApJ...770...57B; 2015MNRAS.446..521S; 2015MNRAS.451.2663H; 2018MNRAS.473.4077P; 2019MNRAS.486.2827D; 2023MNRAS.526.4978S). In previous works, large flux-limited spectroscopic surveys, particularly the Two Degree Field Galaxy Spectroscopic Survey (2dF-GRS; 2001MNRAS.328.1039C), SDSS, and Galaxy and Mass Assembly (GAMA; 2011MNRAS.413..971D), are commonly employed to investigate the local GSMF (2001MNRAS.326..255C; 2008MNRAS.388..945B; 2009MNRAS.398.2177L; 2022MNRAS.513..439D). However, due to depth limitations, accurate and reliable measurements of the local GSMF from spectroscopic surveys remain constrained to relatively high masses. At lower masses, measurements are typically refined to the Local Volume, which lies within or around the Local Void (2008ApJ...676..184T; 2010Natur.465..565P; 2012AJ....144....4M; 2019Ap.....62..293K; 2020A&A...633A..19B), often leading to underestimation of the GSMF. Using the ELUCID simulation (2014ApJ...794...94W; 2016ApJ...831..164W), which was run based on the reconstructed initial density field of the nearby universe, 2019ApJ...872..180C quantified the underestimation of the GSMF by galaxy surveys with various magnitude limits. They identified significant underestimation at the low-mass end in current surveys, starting at stellar mass $10^{8.0}{\rm M}_{\sun}-10^{9.0}{\rm M}_{\sun}$. Therefore, measuring the low-mass end of the GSMF using spectroscopic surveys alone remains highly challenging.
@@
To reach lower stellar masses, we employ the PAC method in combination with deeper photometric surveys, rather than relying solely on spectroscopic data, as have been achieved in 2022ApJ...939..104X. Notably, comparing to the stellar mass limits of $10^{8.2}{\rm M}_{\sun}$ reached in 2022ApJ...939..104X, we enhance various aspects of the previous PAC method and take a significant step into the $10^{6.0}{\rm M}_{\sun}$ frontier with the DESI Y1 Bright Galaxy Survey (BGS) Bright sample and DECaLS photometric sample. This advance is facilitated by leveraging numerous low-redshift tracers available in BGS, enabling accurate $\bar{n}_{2}w_{{\rm{p}}}$ measurements at lower redshifts. Furthermore, we derive the GSMF for blue and red galaxy samples separately, offering deeper insights into the galaxy population. Our results mark significant progress in the measurement of the local GSMF and have the potential to impact our understanding of dark matter and reionization. Specifically, the number density of galaxies with stellar masses around $10^{6.0}{\rm M}_{\sun}$ becomes sensitive to dark matter particle mass (2000ApJ...542..622C; 2001ApJ...556...93B; 2005PhRvD..71f3534V; 2012MNRAS.420.2318L; 2013MNRAS.428..882M; 2016MNRAS.456.4346H; 2017MNRAS.464.4520B) and to the epoch and physics of reionization (1992MNRAS.256P..43E; 2001ARA&A..39...19L; 2008MNRAS.390..920O; 2017MNRAS.465.3913B; 2018ApJ...863..123B). Additionally, our findings can also contribute to advancing the understanding of dwarf galaxy formation.
@@
We introduce the DESI and DECaLS data in Section 2. Section 3 provides a detailed description of the PAC method, including various improvements made to enhance it. The measurements and the resulting GSMF are presented in Section 4. In Section 5, we provide a concise summary. We adopt the cosmology with $\Omega_{\rm m}=0.268$, $\Omega_{\Lambda}=0.732$ and $H_{0}=71{\rm\ km/s/Mpc}$ throughout the paper (2013ApJS..208...19H).
@@ 2.1 DESI Y1 BGS
DESI is a prominent Stage-IV dark energy survey with the goal of obtaining spectra for approximately 40 million extragalactic objects over a five-year period (2013arXiv1308.0847L; 2016arXiv161100036D; 2016arXiv161100037D; 2022AJ....164..207D; 2024AJ....167...62D; 2024AJ....168...58D). Covering more than 14,000 $\mathrm{deg^{2}}$, the survey is conducted using a multi-object fibre-fed spectrograph mounted on the prime focus panel of the 4-meter Mayall Telescope at Kitt Peak National Observatory (2022AJ....164..207D). The spectrometer operates in the wavelength range of 360098003600-9800\,Å and can allocate fibres to 5,000 objects during a visit (2016arXiv161100037D; 2023AJ....165....9S; 2024AJ....168...95M; 2024AJ....168..245P). Additional pipelines supporting the DESI experiment are detailed in 2023AJ....165..144G; redrock2023; fba; 2023AJ....166..259S; 2023AJ....165...50M. The parent catalogue used for DESI target selections is constructed from Data Release 9 of the DESI Legacy Imaging Surveys (2017PASP..129f4101Z; 2019AJ....157..168D). The photometric data contains three optical bands grzgrz from the DECam Legacy Survey (DECaLS; 2019AJ....157..168D), the Dark Energy Survey (DES; 2005astro.ph.10346T), the Beijing-Arizona Sky Survey (BASS; 2017PASP..129f4101Z) and the Mayall $z$-band Legacy Survey (MzLS).
As part of its core observations, DESI is conducting the Bright Galaxy Survey (BGS; 2023AJ....165..253H). BGS spans 14,000 $\mathrm{deg^{2}}$ footprint and includes low-redshift galaxies with $z<0.6$ that can be observed during bright time, when the night sky is approximately $2.5\times$ brighter than nominal dark conditions. BGS provides two galaxy samples: the BGS Bright sample, a $r<19.5$ magnitude-limited sample of approximately 10 million galaxies, and the BGS Faint sample, a fainter $19.5<r<20.175$ sample of approximately 5 million galaxies selected based on surface brightness and colour. The selection and completeness of the BGS samples are detailed in 2023AJ....165..253H.
@@
where $w_{\rm{comp}}$ corrects for fibre-assignment incompleteness, and $w_{\rm zfail}$ adjusts for changes in the relative redshift success rates (2025JCAP...01..125R).
In Figure 1, we present the redshift and stellar mass distributions of the BGS Bright sample within the DECaLS region. We compute their physical properties by applying the SED fitting code CIGALE (2019A&A...622A.103B), utilizing the grzgrz band fluxes from DECaLS and redshifts from DESI spectra. The calculations are based on the stellar population synthesis models by 2003MNRAS.344.1000B, which employ a 2003PASP..115..763C initial mass function and a star formation history characterized by a delayed exponential function $\phi(t)\approx t\exp{(-t/\tau)}$. We consider three metallicities, specifically $Z/Z_{\odot}=0.4,\ 1,\ \text{and}\ 2.5$, where $Z_{\odot}$ represents the metallicity of the Sun. Additionally, we incorporate the 2000ApJ...533..682C extinction law to account for dust reddening, within the range of 0<E(BV)<0.50<E(B-V)<0.5.
[…]
To calculate the correlation functions, we use a random catalogue provided by DESI (2025JCAP...01..125R), which matches the footprint and redshift distribution of the BGS sample used in our analysis.